# When Reasoning Meets Compression: Understanding the Effects of LLMs Compression on Large Reasoning Models

### v2

@@ Abstract
Compression methods, including quantization, distillation, and pruning, improve the computational efficiency of large reasoning models (LRMs). However, existing studies either fail to sufficiently compare all three compression methods on LRMs or lack in-depth interpretation analysis. In this paper, we investigate how the reasoning capabilities of LRMs are compromised during compression, through performance benchmarking and mechanistic interpretation. To uncover the effects of compression on reasoning performance, we benchmark quantized, distilled, and pruned DeepSeek-R1 models on four reasoning datasets (AIME 2024, FOLIO, Temporal Sequences, and MuSiQue). To precisely locate compression effects on model weights, we adapt difference of means and attribution patching techniques, focusing on the activation of every linear component in compressed LRMs, to interpret fine-grained causal relationships between weights and various reasoning capabilities. This fine-grained interpretation addresses a fundamental question of compression: which weights are the most important for reasoning? Overall, we find dynamically quantized 2.51-bit R1 reaches close-to-R1 performance. With empirical verification, we present three main findings that generalize across both Llama and Qwen: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights — a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
@@ 1 Introduction
Large reasoning models (LRMs) such as DeepSeek-R1 (Guo et al., 2025) excel at complex reasoning tasks. However, due to their large sizes, deploying them can be costly and even infeasible for individuals, which hinders AI democratization. Compression methods including quantization, distillation, and pruning reduce computational resources (e.g., GPU memory and disk space). Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022). Current distillation involves black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
@@
However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025; Srivastava et al., 2025; Feng et al., 2025). Although current quantization and pruning methods claim to preserve the performance of general-purpose LLMs, benchmarking both of them on LRMs with more reasoning-intensive datasets helps compare their collapse point. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider distillation effect on knowledge and reasoning (Huang et al., 2024; Agarwal et al., 2024). Another research gap is the lack of interpretability of compression effects on LRMs. It is necessary to interpret how compression methods affect LRMs, as such analysis can reveal existing bottlenecks and provide guidance for future compression research.
@@
Figure 1: An overview of our pipeline. Left: We benchmark compressed R1 variants on various reasoning tasks. Right: By computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight. We empirically verify our findings on weight importance by selectively quantizing or protecting a module to test its importance.
@@
Therefore, due to the lack of compression works on LRMs, we study this fundamental research question: How are the reasoning capabilities of LRMs compressed during compression? We answer it from two perspectives — performance benchmarking and mechanistic interpretation. We first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance. We test dynamic quantization (Daniel Han & team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar & Alistarh, 2023), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) on R1 (or distilled R1). Then, we apply mechanistic interpretability to quantify weight contribution towards four core reasoning capabilities of LRMs: backtracking, uncertainty estimation, example testing, and adding knowledge. By focusing on the activation of every linear component in compressed LRMs, we adapt difference of means (Arditi et al., 2024) to extract steering vectors and attribution patching (Syed et al., 2023) to compute weight importance. Unlike previous analysis (Venhoff et al., 2025) that only measures layer-wise weight contribution, our weight importance scores offer more fine-grained interpretation of weight contribution, addressing the fundamental compression question of locating important weights. By comparing weight importance scores between compressed LRMs and original LRMs, we quantify the effects of distillation and quantization on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects.. Our analysis framework is shown in Figure 1.
@@
With empirical verification, our key findings are summarized below:
@@
  • The mlp.up_proj in the final layer of R1 distilled models emerges as one of the most important model components, addressing a core concern in pruning and quantization literature: identifying critical weights. Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%.
  • Final-layer modules, along with the mlp.gate_proj of R1 distilled Llama and Qwen, are overly compressed by popular quantization methods, highlighting the need for greater attention to preserving their weight precision. A successful protection of only final-layer MLP modules could raise average accuracy by 6.57%, with gains of up to 23.17% over the state-of-the-art quantization.
@@ 2.1 Background
Recent efforts. Recent benchmarking (Liu et al., 2025) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies. As for compressed LRMs, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation. Our interpretation analysis aims to demystify the effects of LLMs compression on LRMs, providing a systematic understanding of existing compressed LRMs.
@@ 2.2 Mechanistic Interpretation
For our interpretation analysis, we target four core reasoning behaviors following an existing work (Venhoff et al., 2025): backtracking, uncertainty estimation, example testing, and adding knowledge. We prompt GPT-4o to locate token sequences of each behavior from the output tokens of our LRMs. To interpret different compression strategies, we adapt difference of means and attribution patching by computing the activation of every linear module in each layer. This allows us to compute the causal relationship between each weight matrix and our target reasoning behaviors.
@@ 2.3 Decoding Compression Effects
We hypothesize that the importance shift should be minimal in the ideal case, as a compressed LRM should remain as close as possible to its original counterpart (the more reasoning-capable model). When visualizing the importance shift from an LRM to its compressed variant (or from a distilled model to its backbone), we only consider decreases in $\mathbf{RI}_{m\ell}^{c}$. By definition, the relative importance of each weight matrix is normalized to sum to one, so any increase in relative importance necessarily compensates for decreases elsewhere. Since it is more informative to track cases where the $\mathbf{RI}_{m\ell}^{c}$ of a more reasoning-capable model decreases (e.g., when the reasoning capability of a weight matrix is diminished), we set all increases in relative importance to zero.
@@ 2.4 Scope
We study three major LLMs compression paradigms, distillation, quantization, and pruning, making our scope comprehensive enough for investigating the effects of diverse compression methods. For distillation, we select four R1 distilled models: DeepSeek-R1-Distill Llama-70B, Qwen-32B, Llama-8B, and Qwen-7B. For quantization, we select 2.51-, 1.73-, and 1.58-bit models by Unsloth [^2] https://huggingface.co/unsloth/DeepSeek-R1-GGUF as the choices of quantized R1 due to their popularity. We also evaluate AWQ Lin et al. (2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models). Specifically, we use all four methods to perform 4-bit quantization, and use GPTQ, GPTAQ, and ANY3 for 3-bit quantization as well, since many AWQ implementations do not support 3-bit. For pruning, we run SparseGPT Frantar & Alistarh (2023) on our two largest distilled models. We run interpretaion analysis on linear modules of all layers within LRMs.
@@ 3.1 Overall Performance
The overall performance of R1 and its compressed variants are in Table 1. We show the performance of pruned R1-Distill-Llama-70B and R1-Distill-Qwen-32B under 50% sparsity in Table 1, as it is the default sparsity level of current works (Zhang et al., 2024a; Sun et al., 2023). Additional analysis of test-time compute and model collapse behavior are in Appendices G and H.
@@
R1-Distill-Llama 70B Distillation & 50% sparse 23.3 71.6 97.6 64.2 (6.7, 13.49)
@@
R1-Distill-Qwen 32B Distillation & 50% sparse 25.6 75.1 97.9 66.2 (2.3, 9.01)
@@
Comparing Compression Strategies. In Table 1, the 2.51-bit R1 achieves the highest average accuracy overall, since it has the smallest compression ratio. Both R1 distilled Llama-70B and Qwen-32B reach close-to-R1 accuracy scores. On MuSiQue, the 2.51-bit R1 also achieves performance close to original R1. Therefore, 2.51-bit R1 has the best overall performance than other compression strategies. Although R1 may be over-parameterized, a compression method with a smaller ratio can still offer advantages over methods with higher compression ratios. In contrast, pruning only 50% of the weights causes significant degradation, rendering the pruned LRMs unusable. Thus, we do not interpret the effect of pruning in the later sections. As for all distillation-only models, Qwen delivers stronger reasoning performance than Llama (Appendix E).
@@ 3.2 Collapse Point
Table 2: Performance of two distilled models under various sparsity levels. We report the one-pass scores for all models in this table.
@@ 3.3 Compression Impact on Knowledge and Reasoning
In Table 2, although Qwen demonstrates stronger reasoning capabilities than Llama, it has significantly lower EM and F1 scores on MuSiQue. Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen puts itself at a disadvantaged position. In other words, models’ parameter count affects knowledge more than reasoning. In addition, we notice pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024. This shows pruning hurts LRMs’ knowledge memorization more than quantization. When a compression method aggressively removes the weights of an LRM, it is expected that the model’s knowledge will be more seriously affected. This phenomenon can also be seen on our dynamically quantized models in Table 1. Since quantization preserves parameter count and our analysis above shows that many quantized models still retain competitive reasoning capability, quantization is recommended on knowledge-intensive tasks.
@@ 4.1 Locating Important Weights
The left part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior. We observe that the final layer houses several most important linear modules across all four behaviors, with the highest value located at up_proj. Therefore, the up_proj in the final layer (32_up) stands out as the most important component.
Interestingly, this finding generalizes to R1 distilled Qwen-7B, as we also observe this up_proj outlier in the final layer of Qwen in Figure 4. Notably, our finding complements a recent analysis (Shao & Wu, 2025), which claims the most important module for reasoning is o_proj. Since identifying important weights is a core research problem of compression methodologies, our finding is valuable for future works.
@@
Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (left) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (right). Each heatmap displays scores of importance (or importance shift) of every module at each layer, providing a fine-grained analysis of weight contributions to the corresponding reasoning capability. On the right, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3. Every cluster of 4 side-by-side heatmaps (including those displayed below) follow the same scaling to show the precise magnitude of each weight module.
Figure 3: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant.
Table 3: Accuracy after selectively quantizing a single component of R1 distilled Llama-8B (e.g., 1_up means to only quantize the up_proj in the first layer) to 3-bit. Ranking of a component is based its $\sum_{c}\mathbf{I}_{m\ell}$, so “2nd col” refers to second place within its column across all four heatmaps (each column consists of 7 linear modules of a layer). “1st overall” means the global highest ranking.

### v3

# arXiv:2504.02010v3[cs.LG] 02 Mar 2026

@@ Abstract
Compression methods, including quantization, distillation, and pruning, improve the computational efficiency of large reasoning models (LRMs). However, existing studies either fail to sufficiently compare all three compression methods on LRMs or lack in-depth interpretation analysis. In this paper, we investigate how the reasoning capabilities of LRMs are compromised during compression, through performance benchmarking and mechanistic interpretation. To uncover the effects of compression on reasoning performance, we benchmark quantized, distilled, and pruned DeepSeek-R1 models on four reasoning datasets (AIME 2024, FOLIO, Temporal Sequences, and MuSiQue). To precisely locate compression effects on model weights, we adapt difference of means and attribution patching techniques, focusing on the activation of every linear component in compressed LRMs, to interpret fine-grained causal relationships between weights and various reasoning capabilities. This fine-grained interpretation addresses a fundamental question of compression: which weights are the most important for reasoning? Overall, we find dynamically quantized 2.51-bit R1 reaches close-to-R1 performance. With empirical verification, we present three main findings that generalize across both R1 and non-R1 LRMs: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights - a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
@@ 1 Introduction
Large reasoning models (LRMs) such as DeepSeek-R1 (Guo et al., 2025) excel at complex reasoning tasks. However, due to their large sizes, deploying them can be costly and even infeasible for individuals, which hinders AI democratization. Compression methods including quantization, distillation, and pruning reduce computational resources (e.g., GPU memory and disk space). Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han and team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022). Current distillation involves black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar and Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
@@
However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025a; Srivastava et al., 2025; Feng et al., 2025). Although current quantization and pruning methods claim to preserve the performance of general-purpose LLMs, benchmarking both of them on LRMs with more reasoning-intensive datasets helps compare their collapse point. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider distillation effect on knowledge and reasoning (Huang et al., 2024; Agarwal et al., 2024). Another research gap is the lack of interpretability of compression effects on LRMs. It is necessary to interpret how compression methods affect LRMs, as such analysis can reveal existing bottlenecks and provide guidance for future compression research.
@@
Figure 1: An overview of our pipeline. On the left, we benchmark compressed R1 variants on various reasoning tasks. On the right, by computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight matrices. We empirically verify our findings on weight importance by selectively quantizing or protecting a module to test its importance.
@@
Therefore, due to the lack of compression works on LRMs, we study this fundamental research question: How are the reasoning capabilities of LRMs compressed during compression? We answer it from two perspectives: performance benchmarking and mechanistic interpretation. In the main text, we first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance. We test dynamic quantization (Daniel Han and team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar and Alistarh, 2023), AlphaPruning (Lu et al., 2024), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi and Johnson, 2025) on R1 (or distilled R1). Then, we apply mechanistic interpretability to quantify weight contribution towards four core reasoning capabilities of LRMs: backtracking, uncertainty estimation, example testing, and adding knowledge. By focusing on the activation of every linear component in compressed LRMs, we adapt difference of means (Arditi et al., 2024) to extract steering vectors and attribution patching (Syed et al., 2023) to compute weight importance. Unlike previous analysis (Venhoff et al., 2025) that only measures layer-wise weight contribution, our weight importance scores offer more fine-grained interpretation of weight contribution, addressing the fundamental compression question of locating important weights. By comparing weight importance between compressed and original LRMs, we quantify the effects of distillation, quantization, and pruning on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects.. Our analysis framework is in Figure 1.
@@
With empirical verification, our key findings are summarized below for better understanding and improving LRMs compression (these findings also generalize to non-R1 families):
@@
  • The mlp.up_proj in the final layer of R1 distilled models emerges as one of the most important model components, addressing a core concern in pruning and quantization literature: identifying critical weights. Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%, which validates its high importance.
  • Final-layer modules, along with the mlp.gate_proj of R1 distilled Llama and Qwen, are overly compressed by popular quantization methods, highlighting the need for greater attention to preserving their weight precision. A successful protection of only final-layer MLP modules could raise average accuracy by 6.57%, with gains of up to 23.17% over the state-of-the-art quantization. This key finding also applies to current pruning methods.
@@ 2.1 Background
Recent efforts. Recent benchmarking (Liu et al., 2025a) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies. As for compressed LRMs, Unsloth (Daniel Han and team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation. Our interpretation analysis aims to demystify the effects of LLMs compression on LRMs, providing a systematic understanding of existing compressed LRMs.
@@ 2.2 Mechanistic Interpretation
For our interpretation analysis, we target four core reasoning behaviors following an existing work (Venhoff et al., 2025): backtracking, uncertainty estimation, example testing, and adding knowledge. We prompt GPT-4o to locate token sequences of each behavior from the output tokens of our LRMs. Our annotation dataset consists of 120 instances drawn from the four benchmark datasets (30 instances from each), so our interpretability analysis spans four different task types and difficulties. Annotation robustness of GPT-4o is demonstrated in Appendix G.
To interpret different compression strategies, we adapt difference of means and attribution patching by computing the activation of every linear module in each layer. This allows us to compute the causal relationship between each weight matrix and our target reasoning behaviors.
@@ 2.3 Decoding Compression Effects
We hypothesize that the importance shift should be minimal in the ideal case, as a compressed LRM should remain as close as possible to its original counterpart (the more reasoning-capable model). When visualizing the importance shift from an LRM to its compressed variant (or from a distilled model to its backbone), we only consider decreases in $\mathbf{RI}_{m\ell}^{c}$. By definition, the relative importance of each weight matrix is normalized to sum to one, so any increase in relative importance necessarily compensates for decreases elsewhere. Since it is more informative to track cases where the $\mathbf{RI}_{m\ell}^{c}$ of a more reasoning-capable model decreases (e.g., when the reasoning capability of a weight matrix is diminished), we set all increases in relative importance to zero. Additional justification of only visualizing the decreases is in Appendix H.
@@ 2.4 Scope
We study three major LLMs compression paradigms, distillation, quantization, and pruning, making our scope comprehensive enough for investigating the effects of diverse compression methods. For distillation, we select four R1 distilled models: DeepSeek-R1-Distill Llama-70B, Qwen-32B, Llama-8B, and Qwen-7B. For quantization, we select 2.51-, 1.73-, and 1.58-bit models by Unsloth [^2] https://huggingface.co/unsloth/DeepSeek-R1-GGUF as the choices of quantized R1 due to their popularity. We also evaluate AWQ Lin et al. (2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi and Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models). Specifically, we use all four methods to perform 4-bit quantization, and use GPTQ, GPTAQ, and ANY3 for 3-bit quantization as well, since many AWQ implementations do not support 3-bit. For pruning, we run SparseGPT Frantar and Alistarh (2023) and AlphaPruning (Lu et al., 2024) on several distilled models. Inside AlphaPruning, SparseGPT is applied to prune the layers after layerwise compression ratios are computed. We run interpretation analysis on linear modules of all layers within LRMs.
@@ 3 Compression Effects on Reasoning Performance
In this section, we examine the effects of compression on reasoning performance. Interestingly, our key findings from the following sections also generalize to non-R1 model families, as elaborated in Appendix J.
@@ 3.1 Overall Performance
The overall performance of R1 and its compressed variants are in Table 1. We show the performance of pruned R1-Distill-Llama-70B and R1-Distill-Qwen-32B under 50% sparsity in Table 1, as it is the default sparsity level of current works (Zhang et al., 2024a; Sun et al., 2023). Additional analysis of test-time compute and model collapse behavior are in Appendices O and P.
@@
R1-Distill-Llama 70B Distillation & 50% SparseGPT 23.3 71.6 97.6 64.2 (6.7, 13.49)
R1-Distill-Llama 70B Distillation & 50% AlphaPruning 26.7 74.2 97.7 66.2 (5.3, 12.39)
@@
R1-Distill-Qwen 32B Distillation & 50% SparseGPT 25.6 75.1 97.9 66.2 (2.3, 9.01)
@@
R1-Distill-Llama 8B Distillation & 30% AlphaPruning 41.1 68.9 82.1 64.1 (0.3, 4.51)
R1-Distill-Llama 8B Distillation & 50% AlphaPruning 6.7 61.7 79.6 49.3 (0.0, 2.95)
@@
Comparing Compression Strategies. In Table 1, the 2.51-bit R1 achieves the highest average accuracy overall, since it has the smallest compression ratio. Both R1 distilled Llama-70B and Qwen-32B reach close-to-R1 accuracy scores. On MuSiQue, the 2.51-bit R1 also achieves performance close to original R1. Therefore, 2.51-bit R1 has the best overall performance than other compression strategies. Although R1 may be over-parameterized, a compression method with a smaller ratio can still offer advantages over methods with higher compression ratios. In contrast, pruning only 50% of the weights causes significant degradation, rendering the pruned LRMs unusable. Thus, we choose to interpret the effect of pruning with greater caution and specify the details in Appendix I. As for all distillation-only models, Qwen delivers stronger reasoning performance than Llama (Appendix E).
@@ 3.2 Collapse Point
Table 2: Performance of two distilled models under various sparsity levels of SparseGPT. We report the one-pass scores for all models in this table.
@@ 3.3 Compression Impact on Knowledge and Reasoning
In Table 2, although Qwen demonstrates stronger reasoning capabilities than Llama, it has significantly lower EM and F1 scores on MuSiQue. Because MuSiQue requires knowledge memorization under the closed-book setting, the smaller parameter count of Qwen puts itself at a disadvantaged position. In other words, models’ parameter count affects knowledge more than reasoning. In addition, we notice that pruned R1-Distill-Llama-70B collapses between 30% and 40% sparsity on MuSiQue, which is even earlier than on AIME 2024. This shows that pruning hurts LRMs’ knowledge memorization more than quantization. When a compression method aggressively removes the weights of an LRM, it is expected that the model’s knowledge will be more seriously affected. This phenomenon can also be seen on our dynamically quantized models in Table 1. Since quantization preserves parameter count and our analysis above shows that many quantized models still retain competitive reasoning capability, quantization is recommended on knowledge-intensive tasks. Additional experiments on knowledge retrieval are specified in Appendix L.
@@ 4.1 Locating Important Weights
The upper part of Figure 2 presents the weight importance of R1 distilled Llama-8B in four heatmaps, each corresponding to a reasoning behavior. We observe that the final layer houses several most important linear modules across all four behaviors, with the highest value located at up_proj. Therefore, the up_proj in the final layer (32_up) stands out as the most important component.
Interestingly, this finding generalizes to R1 distilled Qwen-7B, as we also observe this up_proj outlier in the final layer of Qwen in Figure 4. Notably, our finding complements a recent analysis (Shao and Wu, 2025), which claims the most important module for reasoning is o_proj. Since identifying important weights is a core research problem of compression methodologies, our finding is valuable for future works. Additional diagnosis based on AWQ also reinforces our finding on the last-layer up projection (Appendix M).
@@
Figure 2: $\mathbf{I}_{m\ell}^{c}$ of DeepSeek-R1-Distill-Llama-8B (upper) and change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to Llama-3.1-8B (lower). Each heatmap displays scores of importance (or importance shift) of every module at each layer, providing a fine-grained analysis of weight contributions to the corresponding reasoning capability. For the lower part, increases in $\mathbf{RI}_{m\ell}^{c}$ are set to 0, as they only offset decreases elsewhere as discussed in 2.3. Every cluster of 4 side-by-side heatmaps (including those displayed below) follow the same scaling to show the precise magnitude of each weight module.
Figure 3: Change of $\mathbf{RI}_{m\ell}^{c}$ from DeepSeek-R1-Distill-Llama-8B to its 4-bit AWQ variant. Justification of only showing the decrease of $\mathbf{RI}_{m\ell}^{c}$ in the main content is specified in Appendix H.
Table 3: Accuracy after selectively quantizing a single component of R1 distilled Llama-8B (e.g., 1_up means to only quantize the up_proj in the first layer) to 3-bit. Ranking of a component is based its $\sum_{c}\mathbf{I}_{m\ell}$, so “2nd col” refers to second place within its column across all four heatmaps (each column consists of 7 linear modules of a layer). “1st overall” means the global highest ranking. We provide additional validation on AIME 2024 scores in Appendix N.